Edelman and gronkowski want to play a game of beach volleybal. If the beach ball has a volume of 261L at a temperature of 502k what will the temperature of the balloon be if the volume decreases to 176L?


1
Expert's answer
2021-02-17T07:48:14-0500

"\frac{V_1}{T_1}" = "\frac{V_2}{T_2}"


"\frac{261L}{502K}" = "\frac{176L}{T_2}"


T2 = 338.5K